OpenStudy (anonymous):

Factor completely: x^2 − 4x + 9

OpenStudy (anonymous):

complete the square. you have x^2-4x+4=9+4

OpenStudy (anonymous):

(x-2)^2=13

OpenStudy (anonymous):

x-2 = +-sqrt13 SO, x=2+sqrt13 or x=2-sqrt13

OpenStudy (anonymous):

you get to imaginary solutions 2+or - 1i squareroot of 5